# Iterative Approach to Reconstructing Neural Disparity Fields from Light-field Data

### v1

@@ III Method
Traditional methods for disparity estimation from light field data follow a ”discrete-to-discrete” approach, which fails to overcome the resolution limitations inherent in disparity representation. By contrast, with their continuous nature, neural field implicit representations provide superior expression capabilities compared to traditional discrete disparity representations, as they are not bound by finite resolution constraints, thus enabling robust modeling of complex disparity variations.
@@ III-A Neural disparity field representation
In contrast to traditional implicit representations of 3D shapes, disparity refers to a two-dimensional (2D) geometric manifold in 3D space that is associated with the depth of the scene. It describes the differences between viewpoints instead of directly representing 3D geometric shapes. The implicit continuous parameterization of NDFs allows the modeling of continuous disparities, making it more appropriate for describing continuous variations in disparity within light fields compared to the conventional 2D matrix representations.
The disparity inherent in a scene within a light field is parameterized as a continuous 2D implicit function using neural fields, termed NDFs. This is achieved by approximating the continuous 2D implicit function using a multiresolution grid and the MLPs network $\mathbf{F}_{\mathbf{\Theta}}$ with activation functions, defined as
@@ III-B Inverse problem solving for disparity reconstruction
In the light field, disparity information is shared across various subviews, allowing viewpoint transitions through disparity propagation. In the two-plane four-parameter light field, assuming that occlusion and camera noise are disregarded, the relationship between image $L_{(u_{0},v_{0})}(\mathbf{x})$ at the central viewpoint $(u_{0},v_{0})$ and image $L_{(u_{i},v_{j})}(\mathbf{x})$ at viewpoint $(u_{i},v_{j})$ can be described as follows:
@@
Certain factors such as occlusion and camera noise render Eq.(2) invalid for real-world scenarios. Therefore, this study introduces the propagation mask $m_{(u,v)}(\mathbf{x})$, with $m_{(u,v)}(\mathbf{x})=0$ in cases of occlusion or noise, and $m_{(u,v)}(\mathbf{x})=1$ otherwise. Based on this, the relationship between the central view image $L_{(u_{0},v_{0})}\left(\mathbf{x}\right)$ and the view image $L_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ in the real scenes is adjusted as follows:
@@ III-B1 Forward problem modeling
The light field is regularly sampled on the viewpoint plane, ensuring that the pixels in the central view image are visible in at least half of the viewpoints, resulting in a stronger coupling between the central view and other views. Therefore, this study employs the disparity propagation process from other views to the central view in the light field to model disparity reconstruction. This process effectively utilizes the constraints of regular sampling on the viewpoint plane, while preserving the characteristics of the light field. This provides a robust forward problem model for disparity reconstruction.
To account for occlusion and camera noise in real scenes with a light field, a forward problem model for disparity reconstruction under angular consistency was established as follows:
$L(d(\mathbf{x}))=\frac{\sum_{(u_{i},v_{j})\in\phi}\left(m_{(u_{i},v_{j})}(\mathbf{x})\cdot L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot d(\mathbf{x})\right)\right)}{\sum_{(u_{i},v_{j})\in\phi}m_{(u_{i},v_{j})}(\mathbf{x})}.$ (4)
where $\phi$ is the set of viewpoints.
@@ III-B2 Inverse problem solving
Loss functions for selected viewpoints.
Traditional stereo-matching algorithms often utilize region-based matching windows instead of individual pixels to enhance the robustness of the matching results[39]. To exploit the local structural information of the light field fully, this study utilized a combination of L1 and mean structural similarity (MSSIM) losses [40]. Total variation (TV) [41] regularization is introduced as a loss term to exploit the sparsity of disparities in the gradient domain.
Given the light field $L_{(u,v)}(\mathbf{x})$ and the coordinate $\mathbf{x}$ of the central view, this study aims to determine the optimal parameters $\Theta^{*}$ by minimizing the objective function:
$\begin{split}&{\Theta^{*}}=\underset{\Theta}{\operatorname{arg}\operatorname{min}}\left|L_{(u_{0},v_{0})}\left(\mathbf{x}\right)-L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)\right|+\\ &\alpha\cdot\left(1-\text{MSSIM}\left(L_{(u_{0},v_{0})}\left(\mathbf{x}\right),L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)\right)\right)+\beta\cdot\Psi\left({\hat{d}_{\Theta}\left(\mathbf{x}\right)}\right),\end{split}$ (5)
where $\hat{d}_{\Theta}\left(\mathbf{x}\right)$ represents the disparity predicted at point $\mathbf{x}$ by the network and $L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)$ is computed value using Eq.(4). $\Psi(\mathbf{x})$ denotes the regularization term for the parameterized disparities, where TV regularization is employed. $\alpha$ and $\beta$ are the two hyperparameters.
This study utilized the L1 + $\alpha\cdot\left(\text{MSSIM}\right)$ metric to evaluate the disparity between viewpoints that inadequately convey information and the central view image. This metric, denoted as distance $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$, is defined as follows:
@@
Given that occluded pixels can effectively transmit information from at least half of the viewpoints, this study proposes a simplified strategy for computing an occlusion mask during disparity propagation. Assuming that all viewpoint masks hold a value of 1, signifying that all viewpoints contribute valuable information, this study proceeds by selecting half of the viewpoints based on a comparison of the ray distances $E_{(u_{i},v_{j})}(\mathbf{x})$. It then selects viewpoints $\Phi_{\text{selected}}$ from which the ray distance is smaller than that of the remaining viewpoints; these selected viewpoints are subsequently utilized for conducting backward propagation of rays. In general, this loss function is expressed by
$\mathcal{L}=\sum_{(u_{i},v_{j})\in\Phi_{\text{selected}}}\left(E_{(u_{i},v_{j})}\left(\mathbf{x}\right)\right)+\Psi\left({\hat{d}_{\Theta}\left(\mathbf{x}\right)}\right).$ (7)
This approach effectively avoids the complex computation of propagation masks during disparity propagation, while simultaneously reducing interference from occlusion and noise factors. Therefore, we term this approach the viewpoint selection loss strategy, which improves the accuracy of depth estimation in occluded scenes by giving precedence to the efficacy of information transmission.
@@
To enhance the computational efficiency and expressive power of neural networks, recent research [42, 43, 28, 29, 30] has explored the possibility of the structured storage of latent features. While fully connected networks combined with positional encoding can record data information through weights, they reduce computational efficiency and limit the network’s capacity to represent complex patterns. Previous studies explored different approaches for storing latent features, such as 3D grids [42]. However, this grid structure suffers from resource wastage because a large portion of the internal units remain empty, resulting in the effective utilization of only surface information. To address this issue and reduce memory and training data requirements, prior work has explored the use of tree-like data structures similar to those of Acorn [43]; however, this approach incurs a larger overhead in dynamically maintaining the tree structure during training. Thus, to overcome these limitations, [28] proposed the use of multi-resolution grids for feature storage.
Considering the disparity characteristics, adjacent pixels often have similar values, reflecting the smoothness of scene depth variations. This necessitates the precise capture of subtle changes using the NDF model. Additionally, the distribution of gradients in the disparity exhibits sparsity, with most pixels having gradients close to zero, which is significant only at the scene or object edges. This sparsity implies that the local pixels share the same features, obviating the need to learn unique feature descriptions for each independent pixel, thereby reducing the training time. Thus, this study constructed an NDF network inspired by the approach proposed in [28]. Its structure comprises a multiresolution grid feature module and a MLPs module, as depicted in Fig. 2. The multiresolution grid feature module, a trainable component, encompasses $L$ grids with resolutions linearly increasing from 32 to 128, and is utilized for feature vector storage. The MLPs module consisted of fully connected layers and LeakyReLU activation functions, comprising two hidden layers, each with 256 neurons.
Figure 2: Neural-disparity-field network architecture. The network architecture comprises multiresolution grid-feature and MLPs modules. The multiresolution grid-feature module is a trainable module that stores Hash encoding feature vectors. MLPs module comprises fully connected layers and LeakyReLU activation functions.
@@
    A set of 2D sampling points $\mathbf{x}$ was extracted from the coordinate system of the reference view.
@@
    The multiresolution grid feature module was utilized to extract the corresponding features of the sampling points. These features were input into the MLPs module to predict the corresponding disparity $\hat{d}_{\Theta}\left(\mathbf{x}\right)$.
@@
    The distance $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ is calculated using Eq.(6) and retain those with $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ smaller than the corresponding rays of the other half of the viewpoints.
@@
    The loss of the retained rays is computed using Eq.(7). Subsequently, backward propagation is performed to endow the network with the ability to learn the disparity.
@@ IV Experiments
To comprehensively assess the performance of the proposed method, we conducted comparative experiments on several light field datasets, including the HCI 4D light field dataset [44], 4D Light Field Benchmark (CVIA Konstanz & HCI Heidelberg) [45], and Stanford Lytro dataset [46].
We evaluated the proposed method in synthetic environments of the HCI 4D light field dataset and 4D Light Field Benchmark to verify its fundamental performance and generalization. We then applied it to real Lytro-captured light-field data for a practical disparity estimation assessment. The ability of this method to directly reconstruct disparities from a network at diverse spatial resolutions was also demonstrated. Finally, an ablation study investigated the impact of combined losses, offering deep theoretical insights and practical guidance for algorithm understanding and optimization.
Experiments were conducted on hardware featuring an NVIDIA GeForce RTX 3080 with 10GB memory. The experimental parameter settings in this section are as follows: a grid feature module with six resolutions linearly increasing within the range from 32 to 128, an MSSIM loss window size of 11 pixels, and hyperparameter values $\alpha$ and $\beta$ ranging between $1\text{e}^{-1}$ and $1\text{e}^{1}$, with slight variations for different scenes. During network training, Gaussian noise with intensities ranging from $1\text{e}^{-2}$ to $1\text{e}^{0}$ was introduced.
@@ IV-A1 Complex scene experiments
In this section, disparity estimation is conducted using synthetic light-field datasets, and compared with the MV [47], LF_OCC [10], SPO [12], and AMN [14] algorithms for evaluation. Experimental results are illustrated in Fig. 3, which shows four scenes from the two datasets.
@@
Figure 3: Disparity results for complex synthetic scenes, with odd rows showing estimated disparity maps and even rows displaying profile curves (red line in reference view) comparing actual (red) and algorithm-estimated (blue) disparities. Scenes include from HCI [44] (Buddha and Papillon) and 4D Light Field Benchmark [45] (Pens and Tomb).
@@
In the Buddha scene, the prominent elements include a Buddha statue, slanted wooden planks, stone pillars, and dice. The intricate textures of Buddha statues and stone pillars pose challenges to the robustness of these algorithms. However, as shown in Fig. 3, all of the aforementioned algorithms effectively estimated the disparities of these elements. Slanted wooden planks tested the ability of the algorithms to handle inclined surfaces, where the proposed method outperformed MV, producing smoother disparity predictions on slanted surfaces. Additionally, the black smooth regions on the dice present challenges to the algorithm performance, where the proposed method demonstrates notable advantages over SPO in handling textureless smooth areas. The profile graphs in the second row of Fig. 3 show that the proposed method exhibited minimal oscillations, except for slight oscillations at the boundaries. In the Papillon scene, the primary elements are butterflies and leaves. The butterfly tail features a weakly textured black area and handling such a region is difficult. The proposed method successfully computed disparities in this area compared with MV, SPO, and AMN. Furthermore, for the boundary areas of the leaves, LF_OCC generates considerable noise owing to occlusion, whereas the proposed method exhibits advantages in handling such occluded regions. The Pens scene comprised multiple Pens arranged at various angles and depths. LF_OCC introduces substantial noise at pen boundaries, whereas the proposed method and other comparative algorithms effectively reconstruct the disparities in this scene. The Tomb scene contains several stone statues, and the results reconstructed using the proposed method and SPO exhibit smoothness closer to the ground truth.
This section provides detailed analyses of the Buddha, Papillon, Pens, and Tomb scenes, covering common scenarios of inclined surfaces, smooth areas, occluded regions, and texture structures. Through comparative analysis, the proposed method demonstrated advantages in handling smooth and weakly textured areas. From the profiles, it is evident that the proposed method maintains minimal deviation across the entire curve, although it exhibits oscillations, particularly at the edges of the scene.

### v2

@@ III Method
To ensure generality, the proposed method is demonstrated using the central viewpoint depth and possesses the ability to reconstruct depth information from alternative viewpoints. Additionally, for LFs, there exists a one-to-one correspondence between the scene depth and disparity, which accurately represents the 3D spatial structure of the scene.
@@ III-A Neural disparity field representation
In contrast to traditional implicit representations of 3D shapes, disparity refers to a two-dimensional (2D) geometric manifold in 3D space that is associated with the depth of the scene. It describes the differences between viewpoints instead of directly representing 3D geometric shapes. The implicit continuous parameterization of NDFs allows the modeling of continuous disparities, making it more appropriate for describing continuous variations in disparity within LFs compared to the conventional 2D matrix representations.
The disparity inherent in a scene within an LF is parameterized as a continuous 2D implicit function using neural fields, termed NDFs. This is achieved by approximating the continuous 2D implicit function using a multiresolution grid and the multilayer perceptrons (MLPs) network $\mathbf{F}_{\mathbf{\Theta}}$ with activation functions, defined as
@@ III-B Inverse problem solving for disparity reconstruction
In the LF, disparity information is shared across various subviews, allowing viewpoint transitions through disparity propagation. In the two-plane four-parameter LF, assuming that occlusion and camera noise are disregarded, the relationship between image $L_{(u_{0},v_{0})}(\mathbf{x})$ at the central viewpoint $(u_{0},v_{0})$ and image $L_{(u_{i},v_{j})}(\mathbf{x})$ at viewpoint $(u_{i},v_{j})$ can be described as follows:
@@
Certain factors such as occlusion and camera noise render (2) invalid for real-world scenarios. Therefore, this study introduces the propagation mask $m_{(u,v)}(\mathbf{x})$, with $m_{(u,v)}(\mathbf{x})=0$ in cases of occlusion or noise, and $m_{(u,v)}(\mathbf{x})=1$ otherwise. Based on this, the relationship between the central view image $L_{(u_{0},v_{0})}\left(\mathbf{x}\right)$ and the view image $L_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ in the real scenes is adjusted as follows:
@@ III-B1 Forward problem modeling
The LF is regularly sampled on the viewpoint plane, ensuring that the pixels in the central view image are visible in at least half of the viewpoints, resulting in a stronger coupling between the central view and other views. Therefore, this study employs the disparity propagation process from other views to the central view in the LF to model disparity reconstruction. This process effectively utilizes the constraints of regular sampling on the viewpoint plane, while preserving the characteristics of the LF. This provides a robust forward problem model for disparity reconstruction.
To account for occlusion and camera noise in real scenes with an LF, a forward problem model for disparity reconstruction under angular consistency was established as follows:
$L(d(\mathbf{x}))=\frac{\sum_{(u_{i},v_{j})\in\phi}\left(m_{(u_{i},v_{j})}(\mathbf{x})\cdot L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot d(\mathbf{x})\right)\right)}{\sum_{(u_{i},v_{j})\in\phi}m_{(u_{i},v_{j})}(\mathbf{x})},$ (4)
where $\phi$ denotes the set of viewpoints in the LF data, excluding the central viewpoint.
@@ III-B2 Inverse problem solving
Loss functions
Traditional stereo-matching algorithms often utilize region-based matching windows instead of individual pixels to enhance the robustness of the matching results[44]. To exploit the local structural information of the LF fully, this study utilized a combination of L1 and Mean Structural SIMilarity (MSSIM) losses [45]. Total variation (TV) [46] regularization is introduced as a loss term to exploit the sparsity of disparities in the gradient domain.
This study utilized the L1 + $\alpha\cdot\left(1-\text{MSSIM}\right)$ metric to evaluate the disparity between viewpoints that inadequately convey information and the central view image. This metric, denoted as distance $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$, is defined as follows:
@@
Given that occluded pixels can transmit effective information from at least half of the viewpoints, this study proposes a simplified strategy to calculate disparity propagation masks. Initially, it assumes that the mask values of all viewpoints are set to 1, implying that all viewpoints are capable of providing effective information. Subsequently, by sorting the ray distances $E_{(u_{i},v_{j})}(\mathbf{x})$ of the viewpoints in the set $\phi$, the method selects half of the viewpoints whose ray distance errors are smaller than those of the other half. These selected viewpoints are then used to perform the backward propagation of loss. The calculation method for the propagation mask is expressed as follows:
$m_{(u_{i},v_{j})}(\mathrm{x})=\left\{\begin{array}[]{l}1,E_{(u_{i},v_{j})}(\mathrm{x})\leq E_{\text{median}}(\mathrm{x})\\ 0,\text{ others }\end{array}\right.,$ (6)
where $E_{\text{median}}(\mathrm{x})$ represents the median of the ray distances in the viewpoint set $\phi$.
In general, this loss function is expressed as:
$\begin{split}\mathcal{L}=&\left|L_{(u_{0},v_{0})}\left(\mathbf{x}\right)-L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)\right|+\\ &\alpha\cdot\left(1-\text{MSSIM}\left(L_{(u_{0},v_{0})}\left(\mathbf{x}\right),L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)\right)\right)+\\ &\beta\cdot\Psi\left({\hat{d}_{\Theta}\left(\mathbf{x}\right)}\right),\end{split}$ (7)
where $\hat{d}_{\Theta}\left(\mathbf{x}\right)$ represents the disparity predicted at point $\mathbf{x}$ by the network and $L(\hat{d}_{\Theta}(\mathbf{x}))$ is computed using (4). $\Psi(\hat{d}_{\Theta}(\mathbf{x}))$ denotes the regularization term, where TV regularization is employed. $\alpha$ and $\beta$ are the two hyperparameters.
This method effectively avoids the complex computation of disparity propagation masks while reducing the interference from occlusion noise. This loss function improves the accuracy of depth estimation in occluded scenes by prioritizing the effectiveness of information transfer.
@@
Considering the disparity characteristics, adjacent pixels often have similar values, reflecting the smoothness of scene depth variations. This necessitates the precise capture of subtle changes using the NDF model. Additionally, the distribution of gradients in the disparity exhibits sparsity, with most pixels having gradients close to zero, which is significant only at the scene or object edges. This sparsity implies that the local pixels share the same features, obviating the need to learn unique feature descriptions for each independent pixel, thereby reducing the training time. Thus, this study constructs an NDF network based on the architecture provided in [38]. Its structure comprises a multiresolution grid feature module and an MLPs module, as depicted in Fig. 3. The multiresolution grid feature module, a trainable component, encompasses $L$ grids with resolutions linearly increasing from 32 to 128, and is utilized for feature vector storage. The MLPs module consisted of fully connected layers and LeakyReLU activation functions, comprising two hidden layers, each with 256 neurons.
Figure 3: NDF network architecture. The network consists of a multiresolution grid-feature module, which stores trainable Hash encoding feature vectors, and an MLP module composed of fully connected layers with LeakyReLU activation functions.
@@
    A set of 2D sampling points $\mathbf{x}$ is extracted from the coordinate system of the reference view.
@@
    The multiresolution grid feature module is utilized to extract the corresponding features of the sampling points. These features are input into the MLPs module to predict the corresponding disparity $\hat{d}_{\Theta}\left(\mathbf{x}\right)$.
@@
    The distance $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ is calculated using (5) and is retained for those with $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ smaller than the corresponding rays of the other half of the viewpoints.
@@
    The loss of the retained rays is computed using (7). Subsequently, backward propagation is performed to endow the network with the ability to learn the disparity.
@@ IV Experiments
To comprehensively assess the performance of the proposed method, we conduct experiments on multiple LF datasets, including the HCI 4D LF dataset [47], the 4D LF Benchmark [48], and the UrbanLF-Syn dataset [49]. Additionally, we test on the Stanford Lytro dataset [50], captured using a Lytro LF camera, and the (New) Stanford LF dataset [51], which features a moving camera mounted on a gantry. All comparison methods are publicly available, and we compare the results with the state-of-the-art methods. Following the recommendations of the respective authors, the parameters of each algorithm are set to the optimal values for different datasets and scenes.
Experiments are conducted on an NVIDIA GeForce RTX 3080 with 10GB memory. The experimental settings include a grid feature module with six resolutions from 32 to 128, an MSSIM loss window size of 11 pixels, and hyperparameters $\alpha=0.25$ and $\beta$ ranging from $1\times 10^{-3}$ to $1\times 10^{1}$ with slight variations. Gaussian noise with intensities ranging from $1\times 10^{-2}$ to $1\times 10^{0}$ is introduced during training. The training process is stopped after $8\times 10^{3}$ iterations, processing LF data of size $9\times 9\times 512\times 512$, which takes approximately 30 minutes on the RTX 3080. Using tiny-cuda-nn [52] reduces the runtime to less than 3 minutes.
@@ IV-A1 Complex scene experiments
Figure 4: Disparity results for complex synthetic scenes, with odd rows showing estimated disparity maps and even rows displaying profile curves (red line in reference view) comparing actual (red) and algorithm-estimated (blue) disparities. Scenes include from HCI [47] (Buddha and Papillon) and 4D LF Benchmark [48] (Pens and Tomb).
In this section, disparity estimation is performed using synthetic LF datasets [47, 48], and compared with classical optimization-based methods that do not require a training dataset: POBR [16], which uses superpixel-based regularization; LF_OCC [13], a depth estimation method incorporating occlusion modeling; SPO [15], based on EPIs; and AMN [18], an adaptive matching norm method. The experimental results, including four scenes from the two datasets, are shown in Fig. 4.
Figure 5: Disparity results for complex synthetic scenes from the UrbanLF-Syn dataset [49]. Image34 contains ground truth, and the corresponding MSE error maps are provided. Lower values are better.
@@
In the Buddha scene, the prominent elements include a Buddha statue, slanted wooden planks, stone pillars, and dice. The intricate textures of Buddha statues and stone pillars pose challenges to the robustness of these algorithms. However, as shown in Fig. 4, all of the aforementioned algorithms effectively estimated the disparities of these elements. Slanted wooden planks tested the ability of the algorithms to handle inclined surfaces, where the proposed method outperformed POBR, producing smoother disparity predictions on slanted surfaces. Additionally, the black smooth regions on the dice present challenges to the algorithm performance, where the proposed method demonstrates notable advantages over SPO in handling textureless smooth areas. The profile graphs in the second row of Fig. 4 show that the proposed method exhibited minimal oscillations, except for slight oscillations at the boundaries. In the Papillon scene, the primary elements are butterflies and leaves. The butterfly tail features a weakly textured black area and handling such a region is difficult. The proposed method successfully computed disparities in this area compared with SPO, and AMN. Furthermore, for the boundary areas of the leaves, LF_OCC generates considerable noise owing to occlusion, whereas the proposed method exhibits advantages in handling such occluded regions. The Pens scene comprised multiple Pens arranged at various angles and depths. LF_OCC introduces substantial noise at pen boundaries, whereas the proposed method and other comparative algorithms effectively reconstruct the disparities in this scene. The Tomb scene contains several stone statues, and the results reconstructed using the proposed method and SPO exhibit smoothness closer to the ground truth.
In addition, we compare our method with optimization-based methods, such as POBR and SPO, as well as publicly available supervised methods, including OACC [26], DistgDisp [27], and SubFocal [28], and unsupervised methods like PnPusp, on the UrbanLF-Syn dataset. The experimental results are shown in Fig. 5. In this experiment, the selected scene is a virtual urban environment equipped with multiple light sources to simulate various lighting conditions. A camera array of 81 virtual cameras with identical configurations captures LF data. None of the algorithms in Fig. 5 reconstruct the road in the scene, while our method recovers more details. In comparison, the disparity reconstruction from POBR loses detail and shows blocky artifacts in some regions. SPO is ineffective in certain smooth areas. Our method, with TV constraints, recovers more detail in these regions than SPO. Compared to supervised methods, these methods perform worse than expected on this dataset, while our method shows relative advantages.
Through comparative analysis, the proposed method demonstrates advantages in handling smooth and weakly textured areas.